A ball is thrown straight up from a point 2 m above the ground. The ball reaches a maximum height of 3 m above its starting point and then falls 5 m to the ground. When the ball strikes the ground, what is its displacement from its starting point?
1
Expert's answer
2019-10-07T11:04:24-0400

The displacement from its starting point:


"s=h_f-h_i=0-2=-2\ m."

So, its displacement from its starting point is 2 meters downwards.